if the marginal propensity to consume (MPC) is equal to 0.7, government increases spending by $X, and the GDP increases by $1000. Calculate $X.

A.

$500

B.

$100.

C.

There is not enough information to answer the question.

D.

$ 400.

E.

$300.
Answer #1

Mpc=0.7 then multiplier=1/1-Mpc=1/1-.7=1/0.3=3.33

multiplier=Change in income/change in Govt expenditure

and GDP increases by 1000 then change in G must be 1000/3.333=300

ans is E